OpenStudy (anonymous):

2.8 x10^-4kg of a gas occupies a volume of 224mL at 0°C and 1.013 x10^5Pa. What is the gas? A.CO B.CO2 C.Cl2 D. any gas

OpenStudy (anonymous):

@maheshmeghwal9 , do you mind on working on this together?

OpenStudy (maheshmeghwal9):

no i won't; is this eudiometry?

OpenStudy (maheshmeghwal9):

or stoichiometry?

OpenStudy (anonymous):

I have no clue lol, but it is in part of my practice book for my entrance exam :)

OpenStudy (maheshmeghwal9):

O.O It must use the formula \[PV=nRT \implies PV= {{m }\over {M}}RT.\]

OpenStudy (anonymous):

so this is what I have so far: 0°C=273K, 1.013x10^5=1atm, V=0.224L and mass= 0.00000028g=2.8x10^-7g

OpenStudy (anonymous):

okay then I should use the formula to solve for n? @maheshmeghwal9

OpenStudy (maheshmeghwal9):

no for "M"

OpenStudy (anonymous):

okay so what is do m, and M stand for?

OpenStudy (anonymous):

because n= moles

OpenStudy (maheshmeghwal9):

m= mass u have; i.e. - 0.00000028g=2.8x10^-7g M = molar wt. of substance whose mass u have 0.00000028g=2.8x10^-7g.

OpenStudy (anonymous):

okay thanks, i'll try that then :) so equation is M=mRT/PV

OpenStudy (maheshmeghwal9):

ya of course:)

OpenStudy (maheshmeghwal9):

Plz tell what value for "M" do u get?

OpenStudy (maheshmeghwal9):

when u do this:)

OpenStudy (anonymous):

ok, i'll :)

OpenStudy (maheshmeghwal9):

:)

OpenStudy (anonymous):

@maheshmeghwal9 i got 2.6 :/

OpenStudy (maheshmeghwal9):

so i think the answer is any gas becoz the molar wt. we gt corresponds to \[H_2 = 2. \space \text{And we can take;} \space \space 2.8 \approx 2.\]

OpenStudy (maheshmeghwal9):

sorry 2.8 ->2.6*

OpenStudy (anonymous):

idk, because the answer should be A

OpenStudy (maheshmeghwal9):

i gt A but from a different method:/

OpenStudy (anonymous):

okay, how did you do it? i have to go grocery shopping so if you could just post your method that will be great! thx btw

OpenStudy (maheshmeghwal9):

k! Since 0.224 l of gas contains 2.8 x 10^-4 kg Therefore 22.4 l {I mean one mole} of gas will contain = 0.028 kg. = 28gm & CO 's molar wt. is also 28gm So Answer {A.}

OpenStudy (maheshmeghwal9):

but i don't know why there is no need of pressure & temperature? :/

OpenStudy (maheshmeghwal9):

'l' doesn't mean one mole there. "i mean one mole" is there

OpenStudy (maheshmeghwal9):

@chmvijay Please help:)

OpenStudy (chmvijay):

the answer is CO

OpenStudy (chmvijay):

do u need how

OpenStudy (anonymous):

okay, i kind of understand how you got it @maheshmeghwal9 , @chmvijay do you know how to get the answer?

OpenStudy (chmvijay):

mean i dint get u

OpenStudy (anonymous):

do you know how to answer the question?

OpenStudy (chmvijay):

yaaa same thing what he has given already the equation just put the vaalues u get the molwecular weight u compare that with given gases molecular weight u will find it

OpenStudy (maheshmeghwal9):

@chmvijay My main problem is that i have found 2 answers from 2 different methods.

OpenStudy (maheshmeghwal9):

but which one is correct?

OpenStudy (chmvijay):

which one is other method by doing that how much answer u got

OpenStudy (chmvijay):

tell me the answer first

OpenStudy (maheshmeghwal9):

the answer told by @bronzegoddess is option 'A'

OpenStudy (chmvijay):

yaaaa it is correct only the answer is CO

OpenStudy (maheshmeghwal9):

& my second solution is Since 0.224 l of gas contains 2.8 x 10^-4 kg Therefore 22.4 l {I mean one mole} of gas will contain = 0.028 kg. = 28gm & CO 's molar wt. is also 28gm So Answer {A.} but i don't know why there is no need of pressure & temperature in this method? :/ & my first solution was PV =m/M RT. but i gt option D.

OpenStudy (maheshmeghwal9):

why i m & @bronzegoddess getting option D in PV=m/M RT method?

OpenStudy (chmvijay):